ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 22, 34 Sofronidis, Nikolas Efstathios The Topological Entropy is continuous on a second category set in Diff^{r}(M), whenever M is a compact manifold and r is a nonnegative integer. In this note we answer problem 41 in the list of 50 problems copiled by J. Palis and C. Pugh that reflect the questions discussed at the symposium at the University of Warwick in 1974 (see, for example, A. B. Katok, A conjecture about entropy, Amer. Math. Soc., Transl. (2), vol. 133, 1986).